OpenStudy (anonymous):

Enter the variable that is easiest to solve for in this system of equations. 6y=10x+5 6y=5x+7

OpenStudy (zzr0ck3r):

hmm both are pretty easy

OpenStudy (anonymous):

is it 6x?

OpenStudy (zzr0ck3r):

i guess set them equal to each other and you can solve for x pretty easy

OpenStudy (anonymous):

x=0.4

OpenStudy (anonymous):

how did u get x=0.4

OpenStudy (anonymous):

5x+7=10x+5 7-5=10x-5x 2=5x 2/5=x x=0.4

OpenStudy (zzr0ck3r):

10x + 5 = 5x+7 5x = 2 x = 2/5 = 4/10 = .4

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

ur welcome